# Data-Driven Characterization of Latent Dynamics on Quantum Testbeds

### v1

@@ Abstract
This paper presents a data-driven approach to learn latent dynamics in superconducting quantum computing hardware. To this end, we augment the dynamical equation of quantum systems described by the Lindblad master equation by a parameterized source term that is trained from device data to capture unknown system dynamics, such as environmental interactions and system noise. We consider a structure preserving augmentation that learns and distinguishes unitary from dissipative latent dynamics parameterized by a basis of linear operators, as well as an augmentation given by a nonlinear feed-forward neural network. Numerical results are presented using data from two different quantum processing units (QPU) at LLNL’s Quantum Device and Integration Testbed. We demonstrate that our interpretable, structure preserving models and nonlinear models are able to improve the prediction accuracy of the Lindblad master equation and accurately model the latent dynamics of the QPUs.
@@ 1 Introduction
With the size of quantum systems increasing rapidly, focus has now shifted to quantum error correction and demonstrations of practical applications of quantum computing. Any such application requires accurate control of such a NISQ system that is plagued with sensitivity to a myriad of noise sources [1]. Computing these control pulses requires knowledge of the underlying quantum mechanical model governing the quantum system, which are often only approximately known. While quantum characterization aims to determine the device parameters that define the numerical model to simulate the quantum dynamics, determining the appropriate forms and parameters of quantum master equations for real systems is challenging and typically requires continuous parameter estimation or discrete operator tomography to fit unknown parameters experimentally [2]. Deducing model parameters such as transition frequencies and decoherence times require designing specific protocols, which become increasingly difficult for larger systems. The resulting parameters are estimates of the true underlying device frequencies, and combined with the shortcomings of the quantum dynamical model (i.e. Lindblad master equation) itself, yield only an approximate prediction of the underlying dynamical processes where the modeling errors propagate directly to numerically optimized control pulses.
@@
In order to reduce the modeling errors, we consider a data-driven approach to quantum characterization that augments an approximate dynamical model of the quantum device by an additional source that is trained from measurement data to capture latent quantum dynamics. The trained source term represents a correction term that accounts for any unknown dynamical processes, arising, for example, from latent environmental interactions and noise processes, drift in system frequencies or other Hamiltonian parameters, losses along the control lines of the device, or qubit cross talk. We include the learnable parameters directly into the right-hand-side of the differential equation that describes the known underlying dynamics (Lindblad’s master equation), rather than modeling the entire dynamical system by a machine learning model. This augmented model describes a Universal Differential Equation (UDE) [3], also termed physics-informed Neural Ordinary Differential Equation [4], and employs a machine learning agent as a correction to an underlying (approximate) physical model described by a differential equation that stems from possible prior knowledge of a dynamical system. A similar framework has successfully been applied to learn high-dimensional structural dynamical systems [4], general nonlinear coupled oscillators [5], and to predict back hole dynamics [6]. Those applications have demonstrated that such a ‘gray-box’ learning approach, which combines black-box machine learning models with a white-box approach that approximately describes underlying dynamics yields, good generalization properties, is resilient to noisy data, and requires a smaller training data set. On the other hand, integrating physical knowledge into the learning approach for dynamical systems has most prominently been done via physics-informed neural networks (PINNs) [7, 8, 9, 10]. These approaches typically train a machine learning agent to predict the state of the dynamical system directly, while enforcing the underlying physics by embedding the residual of the governing equations in the loss function. However, these methods typically assume that the exact governing equations are known.
@@
In the quantum domain, machine learning approaches have recently been applied to a variety of applications [11, 12, 13] including Hamiltonian estimation [14, 15, 16], quantum phase estimation [17], quantum error correction [18, 17], quantum control [19, 20, 20], and to learn state discrimination for multi-qubit readout from data [21]. To predict and accelerate quantum dynamical calclulations, [22] provides a benchmark study of 22 supervised machine learning methods comparing their ability to forecast long-time dynamics. The vast majority of these methods can be considered black-box approaches, where a machine learning model (primarily, neural networks) approximates the input-output relation of quantum dynamical processes directly [23, 24, 25, 26]. Deep quantum neural-networks have previously been used to solve Lindblad master equations for open quantum many-body systems [27, 28] and models governing non-Markovian dynamics [29]. Recurrent neural networks (RNNs) were used in [30] to generate an input-output relationship of dissipative mechanisms to time evolution of measurement outcomes. They have also been applied to predict quantum trajectories for transmon qubits under different microwave control sequences [31]. While these use RNNs to model the input-output relation directly, time-derivative learning using Neural ODEs has been applied to learn corrections to the Hamiltonian as well as a recurrent non-Markovian Lindblad operators and applied to synthetically generated data [32]. Such Neural ODEs have also been extended to encode completely-positive trace-preserving (CPTP) maps by rewriting the system Hamiltonian as well as collapse operators and non-Markovian memory kernels as neural-net-like expressions (RNN layers), demonstrating a comparatively smaller training data set is required for accurate model fidelity [33]. Other gray-box approaches that combine black-box neural networks with numerically evolved quantum dynamical processes have demonstrated superior performance over model fitting learning approaches [34], and allow for better interpretability; see for example for the extraction of noise power spectra [35] and [36] for automatic extraction of the effective environment dimensionality as well as eigenfrequencies of the joint system and environment dynamics.
@@
This work considers such a ‘gray-box’ approach where an approximate model of the underlying physics is available and machine-learning models are trained to act as a correction to the approximate dynamical system, thereby capturing and identifying latent dynamics. In particular, we augment Lindblad’s master equation as a baseline dynamical model by source terms that are trained using measured data. We investigate a structure-preserving approach that aims to distinguish between unitary and dissipative dynamics, and allow for interpretability of the learned model. We also consider an augmentation by a neural network model to capture any nonlinear interaction that might be present. We apply the approach to identify latent dynamics of two superconducting quantum processing units (QPUs) in the Quantum Design and Integration Testbed (QuDIT) at Lawrence Livermore National Laboratory (LLNL).
@@
2 Model of Open Quantum Systems: Lindblad Master Equation
@@
where $\rho\in\mathds{C}^{N\times N}$ is the density matrix, $\cdot$$\cdot$] is the commutator, $H\in\mathds{C}^{N\times N}$ is the Hamiltonian governing unitary dynamics and $\mathcal{L}(\rho)$ is a Lindbladian operator that models decoherence. We consider single qubit dynamics, hence $N=2$, however the approach is straightforward to generalize to multiple qubits. When $\mathcal{L}(\rho)=0$, Eq. (1) is called the Liouville-von Neumann (LvN) equation and represents the dynamics of a closed quantum system in density matrix formalism. We target superconducting quantum devices and model the Hamiltonian in the rotational frame as
@@
where $N_{\mathcal{L}}$ are the total number of decoherence operators $\mathcal{L}_{i}$ (also called jump operators), and $\left\{\cdot,\cdot\right\}$ is the anti-commutator. We here consider $N_{\mathcal{L}}=2$ standard Lindblad operators in our baseline model, namely $\mathcal{L}_{1}:=a$ and $\mathcal{L}_{2}:=a^{\dagger}a$, where $\tau_{1}=1/T_{1}$ and $\tau_{2}=1/T_{2}$ are the energy decay and dephasing rates, respectively.
@@ 3 Data-Driven Modeling Through Universal Differential Equations
In the subsequent sections, we introduce a parameterized trainable source term $\mathcal{S}(\rho)$ that acts as a correction to the dynamical equation (1) and that is trained from data:
@@
The form of this source term is introduced below. We refer to the source-free models as base models which captures already identified (approximate) dynamics. Together with the trained source term, (4) then constitutes a universal differential equation, combining known (approximate) dynamics with a data-driven correction term.
@@ 3.1 Learning Structure Preserving Operators
We first consider a structure preserving ansatz for generating and distinguishing unitary from dissipative dynamics. To that end, we consider a splitting of the learnable source term $S(\rho)$ of the form
@@
where the coefficients $\alpha_{j}\in\mathbb{R}$ are trainable parameters and the second term in the summation shifts the energy spectrum such that the ground state energy is 0 (i.e. $\left\langle 0\right|(H+\mathcal{S}_{H})\left|0\right\rangle=0$). Note that for $N=2$, $\Lambda_{i},~i\in[1,N^{2}-1]$ correspond to the Pauli matrices with $\Lambda_{1}:=\sigma_{x}$, $\Lambda_{2}:=\sigma_{y}$ and $\Lambda_{3}:=\sigma_{z}$.
@@
where we take the collapse/jump operators $\widecheck{\Lambda}_{j}$ to be the upper triangular part of the generalized Gell-Mann matrices $\Lambda_{j}$ and $\gamma_{j}\in\mathbb{R}$ denotes a trainable parameter. The decoherence rate for the collapse process described by a jump operator $\widecheck{\Lambda}_{j}$ is $\gamma_{j}$. Since Lindblad’s master equation is invariant under unitary and inhomogenous transformation, the decomposition into the Hamiltonian and dissipative parts is not unique [38, 39, 40, 41] and hence an alternate basis $\widecheck{\Lambda}$ can be defined. We motivate our choice of basis by the fact that in the case of Hermitian jump operators, the purity of the system fulfills $\frac{d}{dt}(\mathrm{Tr}[\rho^{2}])\leq 0$, a condition not satisfied by the experimentally obtained density matrices. Furthermore, it is easy to see that $\mathcal{S}_{\mathcal{L}}(\rho;\widecheck{\Lambda})=\mathcal{S}_{\mathcal{L}}(\rho;i\widecheck{\Lambda})$ and for $N=2$, we see that $\widecheck{\Lambda}_{1}=i\widecheck{\Lambda}_{2}=a$ which represents $T_{1}$ energy decay. Hence, $\widecheck{\Lambda}_{1}$ and $\widecheck{\Lambda}_{3}$ are sufficient to describe the dissipative dynamics for the single qubit case. Despite this, we also include $\widecheck{\Lambda}_{2}$ in our ansatz for ease of implementation and generalization to larger multi-level systems. For a single qubit system, the perturbation to the energy decay rate is given by $\Delta\tau_{1}=\gamma_{1}+\gamma_{2}$. Furthermore, note that $\mathcal{S}_{\mathcal{L}}(\rho;\widecheck{\Lambda}_{3})=4\mathcal{S}_{\mathcal{L}}(\rho;a^{\dagger}a)$  [^1] This is easily seen by constructing the superoperator form of this jump operator. which represents the $T_{2}$ dephasing processes such that the dephasing rate perturbation is $\Delta\tau_{2}=4\gamma_{3}$. The effective and perturbed decoherence times are then given as $T_{i}=(\tau_{i}+\Delta\tau_{i})^{-1}$.
Learning the unitary and dissipative latent dynamics within this structure-preserving ansatz amounts to identifying coefficients $\alpha_{i},\gamma_{i}\in\mathbb{R},i=1,\dots,N^{2}-1$ that yield the highest fidelity when comparing the dynamics of the augmented model (4) to device data. Note that this choice of parameterizing $\mathcal{S}_{H}$ and $\mathcal{S}_{\mathcal{L}}(\rho)$ ensures that (5) is a generator for Completely-Positive Trace-Preserving (CPTP) maps [42], and therefore, enforces physically meaningful quantum state evolution. That is, it preserves the trace and positive semi-definiteness of density matrices.
@@ 3.2 Learning Neural Network Operators
In the second, more general approach for learning quantum dynamical operators, we model the source term as a feed-forward neural network. In this approach, the source $\mathcal{S}(\rho)$ is an $L$-fold composition of network layers $\mathcal{N}_{l}$
@@
where $\mathrm{\boldsymbol{i}}=(j,k)$ is a multi-index, as they allow for easy computation of the expansion coefficients due to their trace orthonormality (i.e. $\mathrm{Tr}\left(\widehat{H}_{\mathrm{\boldsymbol{i}}}\widehat{H}_{\mathrm{\boldsymbol{j}}}\right)=\delta_{\mathrm{\boldsymbol{ij}}}$).
The trainable parameters of this ansatz are the weights and biases $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ of each layer of the neural network. In the absence of the nonlinear activation function (i.e. $\sigma$ is the identity), then the $L$-fold composition reduces to an affine map. In the results presented below, we will investigate both a nonlinear model using the tanhtanh activation function, as well as an affine model for the source term using the identity activation function.
Unlike the structure-preserving ansatz in Section 3.1, the neural-network ansatz does not guarentee a CPTP map. To obtain a valid density matrix, (i.e. one that is positive semi-definite with $\mathrm{Tr}(\rho)=1$) we apply a spectral filter with renormalization. In particular, we construct filtered density matrices as
@@
where $\left\{\Psi_{i}\right\}_{i=1}^{N}$ and $\left\{\mathcal{E}_{i}\right\}_{i=1}^{N}$ are the eigenvectors and the corresponding filtered and renormalized eigenvalues, respectively, $\left\{E_{i}\right\}_{i=1}^{N}$ are the original (unfiltered) eigenvalues, and the filter function $\mathcal{H}(x)$ is the Heaviside step-function. Due to the increased computational cost of the spectral decomposition, we do not apply the filter during the training phase, but only during the deployment phase of the trained model.

### v2

@@ Abstract
This paper presents a data-driven approach to learn latent dynamics in superconducting quantum computing hardware. To this end, we augment the dynamical equation of quantum systems described by the Lindblad master equation with a parameterized source term that is trained from experimental data to capture unknown system dynamics, such as environmental interactions and system noise. We consider a structure preserving augmentation that learns and distinguishes unitary from dissipative latent dynamics parameterized by a basis of linear operators, as well as an augmentation given by a nonlinear feed-forward neural network. Numerical results are presented using data from two different quantum processing units (QPU) at Lawrence Livermore National Laboratory’s Quantum Device and Integration Testbed. We demonstrate that our interpretable, structure preserving, and nonlinear models are able to improve the prediction accuracy of the Lindblad master equation and accurately model the latent dynamics of the QPUs.
@@ 1 Introduction
As the size of quantum systems increases rapidly to be able to perform complex scientific simultaions, focus has now shifted to quantum error correction and demonstrations of practical applications of quantum computing. Any such application requires accurate control of a Noisy Intermediate-scale Quantum (NISQ) system that is plagued with sensitivity to a myriad of noise sources [1]. Computing these control pulses requires knowledge of the underlying quantum mechanical model governing the quantum system. While quantum characterization aims to determine the device parameters that define the numerical model to simulate the quantum dynamics, determining the appropriate forms and parameters of quantum master equations for real systems is challenging and typically requires continuous parameter estimation or discrete operator tomography to fit unknown parameters experimentally [2]. Deducing model parameters such as transition frequencies and decoherence times requires designing specific protocols, which become increasingly difficult for larger systems. The resulting parameters are estimates of the true underlying device frequencies, and combined with the shortcomings of the quantum dynamical model (i.e. Lindblad master equation) itself, yield only an approximate prediction of the underlying dynamical processes where the modeling errors propagate directly to numerically optimized control pulses. More accurate models will yield more robust control pulses, help identify latent noise processes which can greatly inform hardware development and serve as efficient emulators of the quantum devices.
@@
In order to reduce the modeling errors, we consider a data-driven approach to quantum characterization that augments an approximate dynamical model of the quantum device by an additional source that is trained from measurement data to capture latent quantum dynamics. The trained source term represents a correction that accounts for any unknown dynamical processes arising from many sources: latent environmental interactions and noise processes, drift in system frequencies or other Hamiltonian parameters, losses along the control lines of the device, or qubit cross talk. We include the learnable parameters directly into the Lindblad’s master equation that describes the known underlying dynamics, rather than modeling the entire dynamical system by a machine learning model. This augmented model describes a Universal Differential Equation (UDE) [3], also termed physics-informed Neural Ordinary Differential Equation [4]. UDE employs a machine learning agent as a correction to an underlying (approximate) physical model described by a differential equation that stems from possible prior knowledge of a dynamical system. A similar framework has successfully been applied to learn high-dimensional structural dynamical systems [4], general nonlinear coupled oscillators [5], and to predict back hole dynamics [6]. Those applications have demonstrated that such a ‘gray-box’ learning approach, which combines black-box machine learning models with a white-box approach that approximately models known dynamics yields, good generalization properties, is resilient to noisy data, and requires a smaller training data set. On the other hand, integrating physical knowledge into the learning approach for dynamical systems has most prominently been done via physics-informed neural networks (PINNs) [7, 8, 9, 10]. These approaches typically train a machine learning agent to predict the state of the dynamical system directly, while enforcing the underlying physics by embedding the residual of the governing equations in the loss function. However, these methods typically assume that the exact governing equations are known.
@@
In the quantum domain, machine learning approaches have recently been applied to a variety of applications [11, 12, 13] including Hamiltonian estimation [14, 15, 16], quantum phase estimation [17], quantum error correction [18, 17], quantum control [19, 20, 20], and to learn state discrimination for multi-qubit readout from data [21]. To predict and accelerate quantum dynamical calculations, Rodríguez et al.[22] provides a benchmark study of 22 supervised machine learning methods and compare their ability to forecast long-time dynamics. The vast majority of these methods can be considered as black-box approaches, where a machine learning model (primarily, neural networks) approximates the input-output relation of quantum dynamical processes directly [23, 24, 25, 26]. Deep quantum neural-networks have previously been used to solve Lindblad master equations for open quantum many-body systems [27, 28] and models governing non-Markovian dynamics [29]. Recurrent neural networks (RNNs) were used by Leclerc [30] to generate an input-output relationship of dissipative mechanisms to time evolution of measurement outcomes. They have also been applied to predict quantum trajectories for transmon qubits under different microwave control sequences [31]. While these studies use RNNs to model the input-output relation directly, time-derivative learning using Neural ODEs has been demonstrated on synthetic problems to learn corrections to the Hamiltonian as well as a recurrent non-Markovian Lindblad operators[32]. Such Neural ODEs have also been extended to encode completely-positive trace-preserving (CPTP) maps by rewriting the system Hamiltonian, collapse operators, and non-Markovian memory kernels as neural-net-like expressions (RNN layers), demonstrating that a comparatively smaller training data set is required to achieve accurate model fidelity [33]. Other gray-box approaches that combine black-box neural networks with numerically evolved quantum dynamical processes have demonstrated superior performance over model fitting learning approaches [34], and allow for better interpretability; see for example for the extraction of noise power spectra [35] and [36] for automatic extraction of the effective environment dimensionality as well as eigenfrequencies of the joint system and environment dynamics.
@@
This work considers such a ‘gray-box’ approach where we use the Lindblad’s master equation as a baseline dynamical model and augment it with source terms that are trained using measured data. We investigate a structure-preserving approach that aims to distinguish between unitary and dissipative dynamics, and allow for interpretability of the learned model. We also consider an augmentation by a neural network model to capture any nonlinear interaction that might be present. We apply the approach to identify latent dynamics of two superconducting quantum processing units (QPUs) in the Quantum Design and Integration Testbed (QuDIT) at Lawrence Livermore National Laboratory (LLNL).
@@
2 Baseline Model of Open Quantum Systems: Lindblad Master Equation
@@
where $\rho\in\mathds{C}^{N\times N}$ is the density matrix, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the commutator, $H\in\mathds{C}^{N\times N}$ is the Hamiltonian governing unitary dynamics and $\mathcal{L}(\rho)$ is a Lindbladian operator that models decoherence. We consider single qubit dynamics, hence $N=2$, however the approach is straightforward to generalize to multiple qubits. When $\mathcal{L}(\rho)=0$, Eq. (1) is called the Liouville-von Neumann (LvN) equation and represents the dynamics of a closed quantum system in density matrix formalism. We target superconducting quantum devices and model the Hamiltonian in the rotating frame as
@@
[…]
@@ 3 Data-Driven Modeling Through Universal Differential Equations
In the subsequent sections, we introduce a parameterized trainable source term $\mathcal{S}(\rho)$ that acts as a correction to the dynamical equation Eq. (1) and that is trained from data:
@@
The form of this source term is introduced below. We refer to the source-free models as base models which captures already identified (approximate) dynamics. Together with the trained source term, Eq. (4) then constitutes a universal differential equation, combining known and approximate dynamics with a data-driven correction term.
@@ 3.1 Learning Structure Preserving Operators
We first consider a structure preserving ansatz for generating and distinguishing unitary from dissipative dynamics. First, we consider a splitting of the learnable source term $S(\rho)$ of the form
@@
where the coefficients $\alpha_{j}\in\mathbb{R}$ are trainable parameters and the second term in the summation shifts the energy spectrum such that the ground state energy is 0; i.e. $\left\langle 0\right|(H+\mathcal{S}_{H})\left|0\right\rangle=0$. Note that for $N=2$, $\Lambda_{i}$, where $i=1,2,3$, correspond to the Pauli matrices with $\Lambda_{1}:=\sigma_{x}$, $\Lambda_{2}:=\sigma_{y}$ and $\Lambda_{3}:=\sigma_{z}$.
@@
where we take the collapse/jump operators $\widecheck{\Lambda}_{j}$ to be the upper triangular part of the generalized Gell-Mann matrices $\Lambda_{j}$ and $\gamma_{j}\in\mathbb{R}$ denotes a trainable parameter. The decoherence rate for the collapse process described by a jump operator $\widecheck{\Lambda}_{j}$ is $\gamma_{j}$. Since Lindblad’s master equation is invariant under unitary and inhomogenous transformation, the decomposition into the Hamiltonian and dissipative parts is not unique [38, 39, 40, 41]. Hence, an alternate basis $\widecheck{\Lambda}$ can be defined. We motivate our choice of basis by the fact that in the case of Hermitian jump operators, the purity of the system fulfills $\frac{d}{dt}(\mathrm{Tr}[\rho^{2}])\leq 0$, a condition not satisfied by the experimentally obtained density matrices. Furthermore, it is straightforward to see that $\mathcal{S}_{\mathcal{L}}(\rho;\widecheck{\Lambda})=\mathcal{S}_{\mathcal{L}}(\rho;i\widecheck{\Lambda})$ and for $N=2$, we see that $\widecheck{\Lambda}_{1}=i\widecheck{\Lambda}_{2}=a$ which represents $T_{1}$ energy decay. Hence, $\widecheck{\Lambda}_{1}$ and $\widecheck{\Lambda}_{3}$ are sufficient to describe the dissipative dynamics for a single qubit case. Despite this, we also include $\widecheck{\Lambda}_{2}$ in our ansatz for ease of implementation and generalization to larger multi-level systems. For a single qubit system, the perturbation to the energy decay rate is given by $\Delta\tau_{1}=\gamma_{1}+\gamma_{2}$. Furthermore, note that $\mathcal{S}_{\mathcal{L}}(\rho;\widecheck{\Lambda}_{3})=4\mathcal{S}_{\mathcal{L}}(\rho;a^{\dagger}a)$  [^1] This is easily seen by constructing the superoperator form of this jump operator. which represents the $T_{2}$ dephasing processes such that the dephasing rate perturbation is $\Delta\tau_{2}=4\gamma_{3}$. The effective and perturbed decoherence times are then given as $T_{i}=(\tau_{i}+\Delta\tau_{i})^{-1}$.
Learning the unitary and dissipative latent dynamics within this structure-preserving ansatz amounts to identifying coefficients $\alpha_{i},\gamma_{i}\in\mathbb{R},i=1,\dots,N^{2}-1$ that yield the highest fidelity when comparing the dynamics of the augmented model (4) to device data. This choice of parameterizing $\mathcal{S}_{H}$ and $\mathcal{S}_{\mathcal{L}}(\rho)$ ensures that (5) is a generator for Completely-Positive Trace-Preserving (CPTP) maps [42], and therefore, enforces physically meaningful quantum state evolution.
@@ 3.2 Learning Neural Network Operators
A second, more general approach for learning quantum dynamical operators involves modeling the source term as a feed-forward neural network. In this approach, the source $\mathcal{S}(\rho)$ is an $L$-fold composition of network layers $\mathcal{N}_{l}$
@@
where $\mathrm{\boldsymbol{i}}=(j,k)$ is a multi-index, as they allow easy computation of the expansion coefficients due to their trace orthonormality (i.e. $\mathrm{Tr}\left(\widehat{H}_{\mathrm{\boldsymbol{i}}}\widehat{H}_{\mathrm{\boldsymbol{j}}}\right)=\delta_{\mathrm{\boldsymbol{ij}}}$).
The trainable parameters of this ansatz are the weights, $\mathbf{W}_{l}$, and biases, $\mathbf{b}_{l}$, of each layer of the neural network. In the absence of the nonlinear activation function (i.e. $\sigma$ is the identity), then the $L$-fold composition reduces to an affine map. In the results presented below, we will investigate both a nonlinear model using the $tanh$ activation function, as well as an affine model for the source term using the identity activation function.
Unlike the structure-preserving ansatz in Section 3.1, the neural-network ansatz does not guarentee a CPTP map. To obtain a valid density matrix, we apply a spectral filter with renormalization. In particular, we construct filtered density matrices as
@@
where $\left\{\Psi_{i}\right\}_{i=1}^{N}$ and $\left\{E_{i}\right\}_{i=1}^{N}$ are the eigenvectors and the corresponding unfiltered eigenvalues, respectively, $\left\{\mathcal{E}_{i}\right\}_{i=1}^{N}$ are the filtered and renormalized eigenvalues, and the filter function $\mathcal{H}(x)$ is the Heaviside step-function. Due to the increased computational cost of the spectral decomposition, we do not apply the filter during the training phase, but only during the deployment phase of the trained model.